# Galactic bars are already mature at Cosmic Noon: bar strength and flatness at $z\sim 1.5$

### v1

@@ 5.1 The presence of “mature” bars
There are still gaps in our understanding of the formation and evolution of bars at $z>1$, where disks have higher dispersion in comparison to the local Universe (by a factor $\sim 2$ based on ubler19). Simulations have found evidence that bars can still form in such conditions, albeit with weaker strengths (ghosh23). These bars can also survive for $1-2\,\rm Gyr$ if the disk gas fractions are sufficiently low (<0.6<0.6 blandhawthorn23). There is however some uncertainty regarding the source of the initial $m=2$ purturbation. Although bars have been found to form spontaneously in the above mentioned N-body simulations, fragkoudi25 uses a suite of magneto-hydrodynamical cosmological zoom-in (Auriga) simulations to conclude that the intial $m=2$ purturbation is rather a result of mergers. This is also expected to result in bars that are already “saturated” in length and unable to grow further. What this means for the bar profile is unclear however. zheng25 similarly concluded that interaction-driven bars follow an accelerated evolutionary pathway. The only discussion regarding the effect of interaction on the bar profile has been provided in noguchi96, where they suggest a rapid development of a flat profile using simple analytical models.
@@
We do not observe any companions for eight of the galaxies in our sample, or any sign of clear tidal disruption (BK25). Furthermore, whether the bars in our sample are saturated at birth also cannot be verified with our bar length measurements (Sec. 4.2). The values predicted by the Auriga simulations for high-z saturated bars ($4\,\rm kpc$ fragkoudi25) are larger than our bar length estimates. However, our measurement do agree with the bar lengths from other JWST-based studies (e.g., guo25) at $z\sim 1.5$. Such discrepancies between observations and simulations may reflect differences introduced by measurement methods rather than true physical disagreement, as the maximum-ellipticity we have adopted is known to underestimate the bar length (wozniak95; athanassoula02b; laurikainen02; erwin03).
@@
The saturated bars are also expected to be longer than their local counterparts. Hence, more prudent is a comparison of our results to the barred galaxies at $z\sim 0.3-0.85$ in the S4G sample (kim21), which uses the same method of measurement. Our galaxies fall on the lower end of their distribution (Fig. 7 for absolute lengths; Fig. 8 for normalized lengths). While our study is limited by small-number statistics, other JWST-based surveys (leconte24; guo25) contain similar numbers of galaxies per redshift bin, suggesting that our results are still representative of $z\sim 1.5$ galaxies. Hence, the shorter bars at this epoch in comparison to those at $z<1$ does not agree with such bars being saturated at birth. Therefore, we do not find any evidence that majority of the bars in our sample have been heavily influenced by interactions. Therefore, the mature bars in our sample likely have developed as a result of secular evolution.
5.2 The varying strength and impact of $z>1$ bars
@@
Our first estimate of bar strength is the relative rest-frame near-IR bar flux with respect to the disk ($S_{\mathrm{bar}}$; Sec.4.3). We find that the maturity of a bar does not necessarily translate into higher relative surface brightness: flatter (and possibly longer) bars are not always brighter. A flux-based measure, however, does not capture the gravitational torque due to the bar potential, as it ignores bar morphology. Since longer and flatter bars exert larger torques (lee20), we instead use $Q_{\mathrm{b}}$ (Sec.4.4) for a more complete quantification of the bar strength and its impact on the host disk.
@@
Figure 13 shows that $Q_{\mathrm{b}}$ is inversely proportional to axis ratio, as expected for more elongated bars (laurikainen02). Our values follow the relation observed in the local Universe (lee20). While their study finds a small offset between early- and late-type galaxies, this difference is too subtle to make any comparisons to our limited sample. We also observe that our absolute $Q_{\mathrm{b}}$ values at $z\sim 1.5$ lie within the average and lower end of scatter of the $z<0.01$ sample. We note that disks at these epochs are expected to be dynamically hotter and thicker. In our calculations, we adopt scale height range of $0.8\pm 0.3\,\rm kpc$, and this range is propagated into the $Q_{\mathrm{b}}$ uncertainties. Thus the similarity in $Q_{\mathrm{b}}$ values indicate that despite different dynamical conditions, bars at high redshift exert comparable torques on their host disks as those in the local Universe. This result is consistent with kim21, who found little redshift evolution in $Q_{\mathrm{b}}$ between $0.2<z<0.85$. Our values are also in agreement with theirs ($Q_{\mathrm{b}}=0.3-0.25\pm 0.15$).
@@
Given the large intrinsic scatter in $Q_{\mathrm{b}}$ (lee20; kim21), the marginal decrease of $Q_{\mathrm{b}}$ with Sérsic index (Fig. 14) for seven of the galaxies in our sample cannot be considered conclusive. Nonetheless, studies of local galaxies suggest that bars strengthen as they mature (buamgart86; elmegreen89; kim15; lee19; kim21), which is consistent with our findings. More puzzling are two remaining galaxies (IDs: 852 and 1704) that show flat bar profiles but with only upper limits estimates of $Q_{\mathrm{b}}\lesssim 0.10$, therefore falling below the uncertainty threshold (Sec. 4.4). ID 1704 in particular has a distinct rectangular, plane-of-the-disk profile resembling the $4/1$ orbital family (contopoulos88), which typically emerges only in strong bars. We find that the low $Q_{\mathrm{b}}$ values in these galaxies arise from the combination of bar profile and flux. A dominant bulge could also lower $Q_{\mathrm{b}}$ by increasing $\left<F_{\mathrm{R}}\right$> (kim21). But both galaxies have low bulge-to-total ratios (B/T $\sim 0.1$), at the lower end of our sample’s range (0.050.40.05-0.4). Thus, the question remains: how did these mature bars weaken?
@@
Another possibility is the addition of angular momentum from accreted gas (bournaud05). At $z>1$, galaxies experience high gas accretion rates, with the accretion onto the host halos scaling as $(1+z)^{2.25}$ (genel08; goerdt10; dekel13). This could be playing a contributing role since ID: 852 has the second highest gas fraction [^8] The gas mass is determined using metallicity dependent dust-to-gas ratio, where the dust mass is measured using multi-band SED fitting from the COSMOS2020 catalogue (weaver22). Further details will be provided in a follow-up paper ($f_{\mathrm{gas}}=M_{\mathrm{gas}}/(M_{\mathrm{gas}}+M_{\mathrm{\star}})=0.64\pm 0.12$) in our sample. Finally, minor mergers can perturb stellar orbits and weaken bars (martinezvalpuesta17; ghosh21). Such interactions are expected to be more common at $z\gtrsim 1$ (lotz11; mantha18). Once again, ID: 852 is the only galaxy in our sample that has a companion999we derive a photometric redshift of the companion using the software CIGALE and find it to be in agreement with the spectroscopic redshift of the galaxy. Furthermore, the location of the companion is aligned with one of the spiral arms. Hence, we claim that the structure is a physical companion.
5.3 Bars tracing the evolution of high-z disks
@@
The presence of flat mature bars in $z\sim 1.5$ indicates that at least some disks had already undergone significant dynamical settling by this epoch, albeit the exact time period required to establish the flatness is unclear. To estimate a timescale of evolution for the flat bars in our sample, placing the initial formation epoch at $z\sim 2.5-3.5$, we invoke the results of anderson22. They find the shoulders, which give the characteristic flatness of a bar, appear after the bar evolves for $\sim 1-2\,\rm Gyr$ in a secular disk. The models used are purely N-body, without the inclusion of gas and star-formation. The effects of tidal interactions have not been studied. They do, however, study disks with a range of thicknesses up to $\sim 900\,\rm pc$, hence covering the range observed at $z>1$ (lian24). With these caveats in mind, we still emphasize that the detection of such profiles provides a constraint on the conditions of early disks of our Universe at $z\gtrsim 2$.
@@
The detection of bars up to $z\sim 3.5$ with JWST (e.g., guo23; leconte24; guo25) has already challenged the classical expectations of bars being non-existent beyond $z>1$ (sheth12). These disks with high gas fractions (magdis12; genzel15; saintonge16; tacconi18) and high turbulence-driven (ionized) gas velocity dispersion (labbe03; genzel11; wisnioski15; forster18, found to be $2-5\,\times$ higher than that in the local Universe) were expected not to be conducive for bar formation. More recently, however, ubler19 have presented an extensive study of 535 star-forming disk galaxies, where they use the KMOS instrument on the Very Large Telescope (sharples13) to determine the ionized gas velocity dispersion to be $30-45,\rm kms^{-1}$ at $z\sim 1-2.5$. This is $\sim 2\times$ the values observed in the local Universe ($24\pm 5\,\rm kms^{-1}$ in the GHASP sample; epinat10).
@@
Meanwhile, the molecular gas velocity dispersion is found to be $\sim 10-15\,\rm kms^{-1}$ lower, including in galaxies at $z>1$. Hence the molecular gas located in star-forming disks is dynamically decoupled from the low-mass ionized gas components. Meanwhile, the structure of bars and spirals is affected by the molecular and stellar components of the disk (the stars inherit the velocity dispersion of the molecular gas; vanDonkelaar22), rather than the ionized gas (jog84; romeo13). Hence the velocity dispersion directly influencing the formation and evolution of secular structures like bars and spirals would rather be limited to $\sim 30\,\rm kms^{-1}$ at least up to $z\sim 2.5$. Hence, the extreme conditions that were thought to prevent the formation of bars may not be that extreme.
@@
Our results, however, do not simply indicate the presence of bars at $z>1.5$, but also indicate their long-term survival. This conclusion extends the discussion of bars at high-z beyond the emergence of the $m=2$ modes, and into sustained evolution for over $\sim 1-2\,\rm Gyr$. Using the AGAMA/RAMSES hydrodynamical N-body simulations, blandhawthorn24 explore the formation and evolution of bars in high-z turbulent disks. They find that galaxies with $f_{\mathrm{gas}}<0.6$ can have bars forming and evolving for at least $2\,\rm Gyr$. It is hence worth noting that seven out of the nine galaxies in our sample have $f_{\mathrm{gas}}<0.6$. Out of the two exceptions, ID: 852 has a weak bar, as has been discussed in the previous section. The other exception is ID: 1749, which has a strong bar. But one could speculate that the high gas fraction could be due to a more recent accretion of gas and has not influenced the bar for the majority of its lifetime. Nevertheless, larger statistics would be required to establish a firm connection between gas fractions and bar strengths.
@@
Finally, fujii18 provides a theoretical framework on predicting the disk dominance after disk settling101010a disk is considered settled when the stellar and gas components are in equilibrium with the dark matter halo and the stellar bulge, at the epoch of bar formation. blandhawthorn23 verifies the relation using 3D hydro simulations. They conclude that the timescale of bar formation rises exponentially with the decrease in the disk dominance parameter, $f_{\mathrm{disk}}$, which is defined as the square of the ratio of the circular velocities for the disk and the whole galaxy, measured at $2.2\times$ the scale length of the disk (Eq. 11 in fujii18). Once formed, they remain stable and evolve for at least $\sim 1-2\,\rm Gyr$. Hence, the mature bars in our study verify this prediction, since we know these bars are not simply transient $m=2$ perturbations. Furthermore, with the expectation of these bars forming at $z\sim 2.5-3.5$, we can also deduce that the $f_{\mathrm{disk}}$ at that epoch would need to be >0.40.5>0.4-0.5 (based on Fig. 7 in blandhawthorn23). It should be noted that direct measurement of $f_{\mathrm{disk}}$ requires expensive high resolution IFU spectroscopic observations (e.g., price21). Therefore, quantification of the bar profile allows us to reliably determine long-lived (mature) bars which can be used to indirectly estimate parameters like $f_{\mathrm{disk}}$, an estimation of the disk dominance in a galaxy.
6 Summary

### v2

@@ V.1 The presence of “mature” bars
There are still gaps in our understanding of the formation and evolution of bars at $z>1$, where disks have higher dispersion in comparison to the local Universe (by a factor $\sim 2$ based on H. Übler et al., 2019). Simulations have found evidence that bars can still form in such conditions, albeit with weaker strengths (S. Ghosh et al., 2023). These bars can also survive for $1-2\,\rm Gyr$ if the disk gas fractions are sufficiently low (<0.6<0.6 J. Bland-Hawthorn et al., 2023). There is however some uncertainty regarding the source of the initial $m=2$ purturbation. Although bars have been found to form spontaneously in the above mentioned N-body simulations, F. Fragkoudi et al. (2025) uses a suite of magneto-hydrodynamical cosmological zoom-in (Auriga) simulations to conclude that the intial $m=2$ purturbation is rather a result of mergers. This is also expected to result in bars that are already “saturated” in length and unable to grow further. What this means for the bar profile is unclear however. Y. Zheng & J. Shen (2025) similarly concluded that interaction-driven bars follow an accelerated evolutionary pathway. The only discussion regarding the effect of interaction on the bar profile has been provided in M. Noguchi (1996), where they suggest a rapid development of a flat profile using simple analytical models.
@@
We do not observe any companions for eight of the galaxies in our sample, or any sign of clear tidal disruption (BK25). Furthermore, whether the bars in our sample are saturated at birth also cannot be verified with our bar length measurements (Sec. IV.2). The values predicted by the Auriga simulations for high-z saturated bars ($4\,\rm kpc$ F. Fragkoudi et al., 2025) are larger than our bar length estimates. However, our measurement do agree with the bar lengths from other JWST-based studies (e.g., Y. Guo et al., 2025) at $z\sim 1.5$. Such discrepancies between observations and simulations may reflect differences introduced by measurement methods rather than true physical disagreement, as the maximum-ellipticity we have adopted is known to underestimate the bar length (H. Wozniak et al., 1995; E. Athanassoula & A. Misiriotis, 2002; E. Laurikainen & H. Salo, 2002; P. Erwin & L. S. Sparke, 2003).
@@
The saturated bars are also expected to be longer than their local counterparts. Hence, more prudent is a comparison of our results to the barred galaxies at $z\sim 0.3-0.85$ in the S4G sample (T. Kim et al., 2021), which uses the same method of measurement. Our galaxies fall on the lower end of their distribution (Fig. 7 for absolute lengths; Fig. 8 for normalized lengths). While our study is limited by small-number statistics, other JWST-based surveys (Z. A. Le Conte et al., 2024; Y. Guo et al., 2025) contain similar numbers of galaxies per redshift bin, suggesting that our results are still representative of $z\sim 1.5$ galaxies. Hence, the shorter bars at this epoch in comparison to those at $z<1$ does not agree with such bars being saturated at birth. Therefore, we do not find any evidence that majority of the bars in our sample have been heavily influenced by interactions. Therefore, the mature bars in our sample likely have developed as a result of secular evolution.
V.2 The varying strength and impact of $z>1$ bars
@@
Our first estimate of bar strength is the relative rest-frame near-IR bar flux with respect to the disk ($S_{\mathrm{bar}}$; Sec.IV.3). We find that the maturity of a bar does not necessarily translate into higher relative surface brightness: flatter (and possibly longer) bars are not always brighter. A flux-based measure, however, does not capture the gravitational torque due to the bar potential, as it ignores bar morphology. Since longer and flatter bars exert larger torques (Y. H. Lee et al., 2020), we instead use $Q_{\mathrm{b}}$ (Sec.IV.4) for a more complete quantification of the bar strength and its impact on the host disk.
@@
Figure 13 shows that $Q_{\mathrm{b}}$ is inversely proportional to axis ratio, as expected for more elongated bars (E. Laurikainen & H. Salo, 2002). Our values follow the relation observed in the local Universe (Y. H. Lee et al., 2020). While their study finds a small offset between early- and late-type galaxies, this difference is too subtle to make any comparisons to our limited sample. We also observe that our absolute $Q_{\mathrm{b}}$ values at $z\sim 1.5$ lie within the average and lower end of scatter of the $z<0.01$ sample. We note that disks at these epochs are expected to be dynamically hotter and thicker. In our calculations, we adopt scale height range of $0.8\pm 0.3\,\rm kpc$, and this range is propagated into the $Q_{\mathrm{b}}$ uncertainties. Thus the similarity in $Q_{\mathrm{b}}$ values indicate that despite different dynamical conditions, bars at high redshift exert comparable torques on their host disks as those in the local Universe. This result is consistent with T. Kim et al. (2021), who found little redshift evolution in $Q_{\mathrm{b}}$ between $0.2<z<0.85$. Our values are also in agreement with theirs ($Q_{\mathrm{b}}=0.3-0.25\pm 0.15$).
@@
Given the large intrinsic scatter in $Q_{\mathrm{b}}$ (Y. H. Lee et al., 2020; T. Kim et al., 2021), the marginal decrease of $Q_{\mathrm{b}}$ with Sérsic index (Fig. 14) for seven of the galaxies in our sample cannot be considered conclusive. Nonetheless, studies of local galaxies suggest that bars strengthen as they mature (C. W. Baumgart & C. J. Peterson, 1986; B. G. Elmegreen & D. M. Elmegreen, 1989; T. Kim et al., 2015; Y. H. Lee et al., 2019; T. Kim et al., 2021), which is consistent with our findings. More puzzling are two remaining galaxies (IDs: 852 and 1704) that show flat bar profiles but with only upper limits estimates of $Q_{\mathrm{b}}\lesssim 0.10$, therefore falling below the uncertainty threshold (Sec. IV.4). ID 1704 in particular has a distinct rectangular, plane-of-the-disk profile resembling the $4/1$ orbital family (G. Contopoulos, 1988), which typically emerges only in strong bars. We find that the low $Q_{\mathrm{b}}$ values in these galaxies arise from the combination of bar profile and flux. A dominant bulge could also lower $Q_{\mathrm{b}}$ by increasing $\left<F_{\mathrm{R}}\right$> (T. Kim et al., 2021). But both galaxies have low bulge-to-total ratios (B/T $\sim 0.1$), at the lower end of our sample’s range (0.050.40.05-0.4). Thus, the question remains: how did these mature bars weaken?
@@
Another possibility is the addition of angular momentum from accreted gas (F. Bournaud et al., 2005a). At $z>1$, galaxies experience high gas accretion rates, with the accretion onto the host halos scaling as $(1+z)^{2.25}$ (S. Genel et al., 2008; T. Goerdt et al., 2010; A. Dekel et al., 2013). This could be playing a contributing role since ID: 852 has the second highest gas fraction [^8] The gas mass is determined using metallicity dependent dust-to-gas ratio, where the dust mass is measured using multi-band SED fitting from the COSMOS2020 catalogue (J. R. Weaver et al., 2022). Further details will be provided in a follow-up paper ($f_{\mathrm{gas}}=M_{\mathrm{gas}}/(M_{\mathrm{gas}}+M_{\mathrm{\star}})=0.64\pm 0.12$) in our sample. Finally, minor mergers can perturb stellar orbits and weaken bars (I. Martinez-Valpuesta et al., 2017; S. Ghosh et al., 2021). Such interactions are expected to be more common at $z\gtrsim 1$ (J. M. Lotz et al., 2011; K. B. Mantha et al., 2018). Once again, ID: 852 is the only galaxy in our sample that has a companion999we derive a photometric redshift of the companion using the software CIGALE and find it to be in agreement with the spectroscopic redshift of the galaxy. Furthermore, the location of the companion is aligned with one of the spiral arms. Hence, we claim that the structure is a physical companion.
V.3 Bars tracing the evolution of high-z disks
@@
The presence of flat mature bars in $z\sim 1.5$ indicates that at least some disks had already undergone significant dynamical settling by this epoch, albeit the exact time period required to establish the flatness is unclear. To estimate a timescale of evolution for the flat bars in our sample, placing the initial formation epoch at $z\sim 2.5-3.5$, we invoke the results of S. R. Anderson et al. (2022). They find the shoulders, which give the characteristic flatness of a bar, appear after the bar evolves for $\sim 1-2\,\rm Gyr$ in a secular disk. The models used are purely N-body, without the inclusion of gas and star-formation. The effects of tidal interactions have not been studied. They do, however, study disks with a range of thicknesses up to $\sim 900\,\rm pc$, hence covering the range observed at $z>1$ (J. Lian & L. Luo, 2024). With these caveats in mind, we still emphasize that the detection of such profiles provides a constraint on the conditions of early disks of our Universe at $z\gtrsim 2$.
@@
The detection of bars up to $z\sim 3.5$ with JWST (e.g., Y. Guo et al., 2023; Z. A. Le Conte et al., 2024; Y. Guo et al., 2025) has already challenged the classical expectations of bars being non-existent beyond $z>1$ (K. Sheth et al., 2012). More recently, C. de Sá-Freitas et al. (2025) have also found that bars in the local Universe might have existed well before $z\sim 1$ based on stellar age estimates. These disks with high gas fractions (G. E. Magdis et al., 2012; R. Genzel et al., 2015; A. Saintonge et al., 2016; L. J. Tacconi et al., 2018) and high turbulence-driven (ionized) gas velocity dispersion (I. Labbé et al., 2003; R. Genzel et al., 2011; E. Wisnioski et al., 2015; N. M. Förster Schreiber et al., 2018, found to be $2-5\,\times$ higher than that in the local Universe) were expected not to be conducive for bar formation. More recently, however, H. Übler et al. (2019) have presented an extensive study of 535 star-forming disk galaxies, where they use the KMOS instrument on the Very Large Telescope (R. Sharples et al., 2013) to determine the ionized gas velocity dispersion to be $30-45,\rm kms^{-1}$ at $z\sim 1-2.5$. This is $\sim 2\times$ the values observed in the local Universe ($24\pm 5\,\rm kms^{-1}$ in the GHASP sample; B. Epinat et al., 2010).
@@
Meanwhile, the molecular gas velocity dispersion is found to be $\sim 10-15\,\rm kms^{-1}$ lower, including in galaxies at $z>1$. Hence the molecular gas located in star-forming disks is dynamically decoupled from the low-mass ionized gas components. Meanwhile, the structure of bars and spirals is affected by the molecular and stellar components of the disk (the stars inherit the velocity dispersion of the molecular gas; F. van Donkelaar et al., 2022), rather than the ionized gas (C. J. Jog & P. M. Solomon, 1984; A. B. Romeo & N. Falstad, 2013). Hence the velocity dispersion directly influencing the formation and evolution of secular structures like bars and spirals would rather be limited to $\sim 30\,\rm kms^{-1}$ at least up to $z\sim 2.5$. Hence, the extreme conditions that were thought to prevent the formation of bars may not be that extreme.
@@
Our results, however, do not simply indicate the presence of bars at $z>1.5$, but also indicate their long-term survival. This conclusion extends the discussion of bars at high-z beyond the emergence of the $m=2$ modes, and into sustained evolution for over $\sim 1-2\,\rm Gyr$. Using the AGAMA/RAMSES hydrodynamical N-body simulations, J. Bland-Hawthorn et al. (2024) explore the formation and evolution of bars in high-z turbulent disks. They find that galaxies with $f_{\mathrm{gas}}<0.6$ can have bars forming and evolving for at least $2\,\rm Gyr$. It is hence worth noting that seven out of the nine galaxies in our sample have $f_{\mathrm{gas}}<0.6$. Out of the two exceptions, ID: 852 has a weak bar, as has been discussed in the previous section. The other exception is ID: 1749, which has a strong bar. But one could speculate that the high gas fraction could be due to a more recent accretion of gas and has not influenced the bar for the majority of its lifetime. Nevertheless, larger statistics would be required to establish a firm connection between gas fractions and bar strengths.
@@
Finally, M. S. Fujii et al. (2018) provides a theoretical framework on predicting the disk dominance after disk settling101010a disk is considered settled when the stellar and gas components are in equilibrium with the dark matter halo and the stellar bulge, at the epoch of bar formation. J. Bland-Hawthorn et al. (2023) verifies the relation using 3D hydro simulations. They conclude that the timescale of bar formation rises exponentially with the decrease in the disk dominance parameter, $f_{\mathrm{disk}}$, which is defined as the square of the ratio of the circular velocities for the disk and the whole galaxy, measured at $2.2\times$ the scale length of the disk (Eq. 11 in M. S. Fujii et al., 2018). Once formed, they remain stable and evolve for at least $\sim 1-2\,\rm Gyr$. Hence, the mature bars in our study verify this prediction, since we know these bars are not simply transient $m=2$ perturbations. Furthermore, with the expectation of these bars forming at $z\sim 2.5-3.5$, we can also deduce that the $f_{\mathrm{disk}}$ at that epoch would need to be >0.40.5>0.4-0.5 (based on Fig. 7 in J. Bland-Hawthorn et al., 2023). It should be noted that direct measurement of $f_{\mathrm{disk}}$ requires expensive high resolution IFU spectroscopic observations (e.g., S. H. Price et al., 2021). Therefore, quantification of the bar profile allows us to reliably determine long-lived (mature) bars which can be used to indirectly estimate parameters like $f_{\mathrm{disk}}$, an estimation of the disk dominance in a galaxy.
VI Summary